OpenStudy (anonymous):

Don put $3,000 in a savings account with an interest rate of 3% for three years. If the interest is compounded annually, how much money will he have at the end of the three years? A. $3,378.18 B. $6,591.00 C. $3,182.70 D. $3,278.18

OpenStudy (anonymous):

(3000)(1.03)(1.03)(1.03) = (3000)(1.03)^3 = (3000)(1.092727)

OpenStudy (anonymous):

wich one is it

OpenStudy (anonymous):

The "1.03" factors come into play because it is "1" plus "3%" because you have the principal plus the interest after each year. To get the answer, all you have to do is do that last multiplication. Don't worry, I checked, and the answer will be among the choices.

OpenStudy (anonymous):

is it D

OpenStudy (anonymous):

Yes! Good job!

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @Cookie_009